# On the asymptotic profile of solutions to semilinear damped wave equations with critical nonlinearities

### v1

@@ 4.1. Upper bound estimate of lifespan
For all $R$ satisfies $1\leq R_{0}\leq R\leq T_{\varepsilon}$, we define the following functions:
[…]
We can easily see that the support of $\psi_{R}$ is contained in
$Q_{R}:=[0,R]\times B_{\sqrt{R}}\quad\text{ with }\quad B_{\sqrt{R}}:=\left\{x\in\mathbb{R}^{n},\,\,|x|\leq\sqrt{R}\right\}.$
Next, we define the quantities
$I_{R}$ $:=\int_{Q_{R}}\mathcal{N}(u(t,x))\psi_{R}(t,x)d(t,x),$
$y=y(r)$ $:=\int_{Q_{r}}\mathcal{N}(u(t,x))\psi^{*}_{r}(t,x)d(t,x)$
and
$Y(R):=\int_{0}^{R}y(r)r^{-1}dr.$
Multiplying both sides of the main equation in (1) by the test function $\psi_{R}(t,x)$ and performing integration by parts over $0,\infty)\times\mathbb{R}^{n}$, we get
$I_{R}+\varepsilon\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))\psi_{R}(0,x)dx=\int_{Q_{R}}u(t,x)\left(\partial_{t}^{2}-\Delta-\partial_{t}\right)\psi_{R}(t,x)d(t,x).$
Using the dominated convergence theorem combined with the condition (11), there exists a constant $R_{0}\geq 1$ satisfies
$\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))\psi_{R}(0,x)dx>\frac{1}{2}\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))dx>0,$ (31)
for all $R\geq R_{0}$. Without loss of generality, we may assume that $T_{\varepsilon}\geq R_{0}$. In [8, Theorem 5], the authors proved that
$\frac{Y(R)}{\log\left(2\right)}+\varepsilon\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))\psi_{R}(0,x)dx$ $\leq I_{R}+\varepsilon\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))\psi_{R}(0,x)dx$
$\leq C_{1}R^{\frac{n}{2}}\mathcal{N}^{-1}\left(\frac{Y^{\prime}(R)}{C_{0}R^{\frac{n}{2}}}\right),$ (32)
for all $R\geq R_{0}$. As a direct consequence, the relations (31) and (32) yield
$\mathcal{N}\left(C_{2}(\varepsilon+Y(R))R^{-\frac{n}{2}}\right)\leq C_{0}R^{-\frac{\mathrm{n}}{2}}Y^{\prime}(R)\text{ for any }R_{0}\leq R\leq T_{\varepsilon}$
and
$\mathcal{N}\left(C_{3}R^{-\frac{n}{2}}Y(R)\right)\leq C_{0}R^{-\frac{n}{2}}Y^{\prime}(R)\text{ for any }R_{0}\leq R\leq T_{\varepsilon}.$
The first estimate deduces
$Y^{\prime}(R)\geq C_{4}(\varepsilon+Y(R_{0}))^{1+\frac{2}{n}}R^{-1}\mu\left(C_{2}(\varepsilon+Y(R_{0}))R^{-\frac{n}{2}}\right)\gtrsim\varepsilon^{1+\frac{2}{n}}R^{-1}\mu\left(C_{5}R^{-\frac{n}{2}}\right)$ (33)
and the second deduces
$R^{-1}\mu\left(C_{5}R^{-\frac{n}{2}}\right)\lesssim\frac{Y^{\prime}(R)}{Y(R)^{1+\frac{2}{n}}}.$ (34)
We introduce the following functions
$\Psi_{R_{0}}(s)=\int_{R_{0}}^{s}\frac{\mu\left(Cr^{-\frac{n}{2}}\right)}{r}\;dr\text{ and }\phi(s):=\Psi_{R_{0}}^{-1}\left(\frac{\Psi_{R_{0}}(s)}{2}\right)\text{ for }s\geq R_{0}.$
The second definition implies
$\Psi_{R_{0}}(R)=2\Psi_{R_{0}}(\phi(R)),\quad\text{ that is, }\quad\frac{\Psi_{R_{0}}(R)}{2}=\int\limits_{R_{0}}^{\phi(R)}\frac{\mu\left(Cr^{-\frac{n}{2}}\right)}{r}dr=\int\limits_{\phi(R)}^{R}\frac{\mu\left(Cr^{-\frac{n}{2}}\right)}{r}dr.$
After denoting $r:=R$ and integrating two sides of (33) over $\left[R_{0},\phi(R)\right]$, one finds
$\varepsilon^{1+\frac{2}{n}}\int\limits_{R_{0}}^{\phi(R)}\frac{\mu\left(Cr^{-\frac{n}{2}}\right)}{r}dr\lesssim Y(\phi(R))-Y\left(R_{0}\right)\lesssim Y(\phi(R)).$ (35)
In the same way, denoting $r:=R$ and integrating two sides of (34) over $\phi(R),R]$, one has
$\int\limits_{\phi(R)}^{R}\frac{\mu\left(Cr^{-\frac{n}{2}}\right)}{r}dr\lesssim\frac{n}{2}Y(\phi(R))^{-\frac{2}{n}}.$ (36)
Combining (35) and (36), we obtain
$\Psi_{R_{0}}(R)\lesssim\varepsilon^{-\frac{2}{n}\left(1+\frac{2}{n}\right)}(\Psi_{R_{0}}(R))^{-\frac{2}{n}},\text{ that is, }\quad\Psi_{R_{0}}(R)\lesssim\varepsilon^{-\frac{2}{n}}.$
Therefore, we can easily see that
$\Psi(R)=\Psi_{R_{0}}(R)+\int_{1}^{R_{0}}\frac{\mu\left(Cr^{-\frac{n}{2}}\right)}{r}\;dr\lesssim\varepsilon^{-\frac{2}{n}}.$
Finally, letting $R\rightarrow T_{\varepsilon}$ in the last inequality one arrives at
$T_{\varepsilon}\lesssim\Psi^{-1}\left(C\varepsilon^{-\frac{2}{n}}\right),$
which we wanted to prove.
4.2. Lower bound estimate of lifespan
In this subsection, we establish the lower bound estimate of the lifespan
$T_{\varepsilon}\gtrsim\Psi\left(c\varepsilon^{-\frac{2}{n}}\right),$
under the conditions stated in Theorem 1.3. The proof relies on the argument introduced in [13], together with the solution spaces and the quantities established in the proof of Theorem 1.1. Specifically, we recall some estimates related to the global existence of small data solutions.
@@ 4.2. Lower bound estimate of lifespan
Afterwards, motivated by the approach in [13], we determine
@@
Proof of Theorem 1.3. Combining the results from subsections 4.1 and 4.2, we conclude the proof of Theorem 1.3.
@@ Acknowledgment.
This work is supported by Vietnam Ministry of Education and Training and Vietnam Institute for Advanced Study in Mathematics under grant number B2026-CTT-04.

### v2

@@ 4. Some remarks to sharp lifespan estimates
$\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))dx>0.$
$T_{\varepsilon}\lesssim\Psi^{-1}\left(C\varepsilon^{-\frac{2}{n}}\right)$ (34)
satisfies the lower bound estimate
$T_{\varepsilon}\gtrsim\Psi^{-1}\left(c\varepsilon^{-\frac{2}{n}}\right),$ (35)
Remark 4.1.
Linking the achieved estimates in Proposition 4.1, one recognizes that the lifespan $T_{\varepsilon}$ in spatial dimensions $1\leq n\leq 4$ under the non-Dini condition are determined by the following relations:
$\Psi(T_{\varepsilon})\sim\varepsilon^{-\frac{2}{n}}.$
In fact, under the assumptions of Proposition 4.1, the authors in [4] proved (34) for all $n\geq 1$ and (35) for $n=1$. In addition, the approach in [12] can be easily applied to derive an estimate (35) for $n=2,3$. Therefore, the novelty of Proposition 4.1 lies in establishing estimate (35) for $n=4$.
Proof of the estimate (35). The proof relies on the argument introduced in [14], together with the solution spaces and the quantities established in the proof of Theorem 1.1. Specifically, we recall some estimates related to the global existence of small data solutions.
@@
Afterwards, we consider
@@ Acknowledgment.
This work is supported by Vietnam Ministry of Education and Training and Vietnam Institute for Advanced Study in Mathematics under grant number B2026-CTT-04. The authors would like to thank Prof. Tuan Anh Dao (Hanoi University of Science and Technology), Prof. Wenhui Chen (Guangzhou University), and Prof. Giovanni Girardi (Università Politecnica delle Marche) for their helpful advice in the preparation of this paper.